# On semipositone problems over ${\mathbb{R}}^{N}$ for the fractional $p$-Laplace operator

### v3

@@ 3. Existence and qualitative properties of the mountain pass solutions
Since $u_{a_{n}}\rightarrow\tilde{u}$ in $L^{\infty}({\mathbb{R}}^{N})$ (Proposition 3.6) and $\tilde{u}>0$ in ${\mathbb{R}}^{N}$ (Proposition 3.8), there exists $\eta_{1}$ depending on $R_{1}$ and $n_{2}\in{\mathbb{N}}$ such that $u_{a_{n}}\geq\eta_{1}$ on $\overline{B_{R_{1}}}$ for all $n\geq n_{2}$. For $10<\varepsilon<$ 1 as in (3.41) we further choose $\varepsilon<\eta_{1}$. Hence
@@
Acknowledgements: The first author is partly funded by the Department of Atomic Energy, Government of India, under project no. 12-R&\&&D-TFR-5.01-0520. The second author acknowledges the support of the CSIR fellowship, file no. 09/1125(0016)/2020–EMR–I. The authors express gratitude to Prof. Abhishek Sarkar (IIT Jodhpur) for beneficial discussions.
@@
$=C(\beta,p)\bigg{|}(a^{-})_{M}^{\frac{\beta+p-1}{p}}-(b^{-})_{M}^{\frac{\beta+p-1}{p}}\bigg{|}^{p}$
@@
$l_{0}=p_{s}^{*},\quad l_{j+1}=\gamma l_{j}+\frac{p^{*}_{s}(p-1)}{p},\text{ and }\gamma=\frac{p^{*}_{s}}{p\bar{r}^{\prime}},$ (A.2)
where $\bar{r}^{\prime}=\frac{\bar{r}}{\bar{r}-1}$ and $\bar{r}$ is given by (3.23). Then there exist $\beta_{1},\beta_{2}>0$ such that $\beta_{1}<\frac{\gamma^{j}}{l_{j}}<\beta_{2}$ for all $j\in{\mathbb{N}}$.
@@
$l_{j}=\gamma^{j}l_{0}+\frac{p_{s}^{*}(p-1)}{p}\sum\limits_{i=0}^{j-1}\gamma^{i}=\gamma^{j}p_{s}^{*}+\frac{p_{s}^{*}(p-1)}{p}\sum\limits_{i=0}^{j-1}\gamma^{i}.$ (A.3)
Dividing (A.3) by $\gamma^{j}$, we obtain
$\frac{l_{j}}{\gamma^{j}}$ $=p_{s}^{*}+\frac{p_{s}^{*}(p-1)}{p}\left(\frac{1}{\gamma^{j}}+\frac{1}{\gamma^{j-1}}+\cdots+\frac{1}{\gamma}\right)=p_{s}^{*}+\frac{p_{s}^{*}(p-1)}{p}\sum\limits_{i=1}^{j}\frac{1}{\gamma^{i}}<p_{s}^{*}+\frac{p_{s}^{*}(p-1)}{p}\sum\limits_{i=1}^{\infty}\frac{1}{\gamma^{i}}.$
@@
$\frac{l_{j}}{\gamma^{j}}<p_{s}^{*}+\frac{p_{s}^{*}(p-1)}{p}\frac{1}{\gamma-1}.$ (A.4)
@@
$\frac{l_{j}}{\gamma^{j}}>p_{s}^{*}.$ (A.5)
The combination of (A.4) and (A.5) yields $\beta_{1}<\frac{\gamma^{j}}{l_{j}}<\beta_{2},$ for all $j\in{\mathbb{N}}$, where $\beta_{1}=\frac{1}{p_{s}^{*}+\frac{p_{s}^{*}(p-1)}{p(\gamma-1)}}$ and $\beta_{2}=\frac{1}{p_{s}^{*}}$. ∎

### v4

@@ 3. Existence and qualitative properties of the mountain pass solutions
Since $u_{a_{n}}\rightarrow\tilde{u}$ in $L^{\infty}({\mathbb{R}}^{N})$ (Proposition 3.6) and $\tilde{u}>0$ in ${\mathbb{R}}^{N}$ (Proposition 3.8), there exists $\eta_{1}>0$ depending on $R_{1}$ and $n_{2}\in{\mathbb{N}}$ such that $u_{a_{n}}\geq\eta_{1}$ on $\overline{B_{R_{1}}}$ for all $n\geq n_{2}$. For $10<\varepsilon<$ 1 as in (3.41) we further choose $\varepsilon<\eta_{1}$. Hence
@@
Acknowledgements: N. Biswas is funded by the Department of Atomic Energy, Government of India, under project no. 12-R&\&&D-TFR-5.01-0520 and the Science and Engineering Research Board, Government of India, through the National Postdoctoral Fellowship (PDF/2023/000038). R. Kumar acknowledges the support of the CSIR fellowship, file no. 09/1125(0016)/2020–EMR–I. The authors thank Prof. Abhishek Sarkar (IIT Jodhpur) for the beneficial discussions.
@@
$=C(\beta,p)\bigg{|}(a^{-})_{M}^{\frac{\beta+p-1}{p}}-(b^{-})_{M}^{\frac{\beta+p-1}{p}}\bigg{|}^{p}.$
@@
$l_{0}=p_{s}^{*},\quad l_{j+1}=\theta l_{j}+\frac{p^{*}_{s}(p-1)}{p},\text{ and }\theta=\frac{p^{*}_{s}}{p\bar{r}^{\prime}},$ (A.2)
where $\bar{r}^{\prime}=\frac{\bar{r}}{\bar{r}-1}$ and $\bar{r}$ is given by (3.23). Then there exist $\beta_{1},\beta_{2}>0$ such that $\beta_{1}<\frac{\theta^{j}}{l_{j}}<\beta_{2}$ for all $j\in{\mathbb{N}}$.
@@
$l_{j}=\theta^{j}l_{0}+\frac{p_{s}^{*}(p-1)}{p}\sum\limits_{i=0}^{j-1}\theta^{i}=\theta^{j}p_{s}^{*}+\frac{p_{s}^{*}(p-1)}{p}\sum\limits_{i=0}^{j-1}\theta^{i}.$ (A.3)
Dividing (A.3) by $\theta^{j}$, we obtain
$\frac{l_{j}}{\theta^{j}}$ $=p_{s}^{*}+\frac{p_{s}^{*}(p-1)}{p}\left(\frac{1}{\theta^{j}}+\frac{1}{\theta^{j-1}}+\cdots+\frac{1}{\theta}\right)=p_{s}^{*}+\frac{p_{s}^{*}(p-1)}{p}\sum\limits_{i=1}^{j}\frac{1}{\theta^{i}}<p_{s}^{*}+\frac{p_{s}^{*}(p-1)}{p}\sum\limits_{i=1}^{\infty}\frac{1}{\theta^{i}}.$
@@
$\frac{l_{j}}{\theta^{j}}<p_{s}^{*}+\frac{p_{s}^{*}(p-1)}{p}\frac{1}{\theta-1}.$ (A.4)
@@
$\frac{l_{j}}{\theta^{j}}>p_{s}^{*}.$ (A.5)
The combination of (A.4) and (A.5) yields $\beta_{1}<\frac{\theta^{j}}{l_{j}}<\beta_{2},$ for all $j\in{\mathbb{N}}$, where $\beta_{1}=\frac{1}{p_{s}^{*}+\frac{p_{s}^{*}(p-1)}{p(\theta-1)}}$ and $\beta_{2}=\frac{1}{p_{s}^{*}}$. ∎